Circular No. 7191

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 W. Cui and L. Wen, Massachusetts Institute of Technology; S. N.

Zhang and X.-B. Wu, University of Alabama, Huntsville, report: “We have detected a periodic signal in the x-ray intensity of blackhole candidate XTE J1550–564, using public data from the All-Sky Monitor (ASM) aboard RXTE. A Lomb-Scargle periodogram is constructed from the ASM light curve of this transient source for the period when it was in a recent x-ray outburst (1998 Sept. 4–1999 May 22). A peak is apparent at 1.398 ± 0.003 days in the periodogram, with a power of ∼ 40 (the Poisson noise power is normalized to 1 in the same units). Taking into account the intrinsic red noise of the source, we estimate that the detection bears a false-alarm probability of ∼ 10^{−4}. We have also tried to break up the outburst light curve into segments and have performed the same analysis on each segment. We have found that the signal is always present at the same position in the periodogram (within measurement uncertainties). Similar results are obtained from Fourier analyses of the light curve. The epoch-folded light curve is roughly of sinusoidal form, with a mean peak-to-peak amplitude ∼ 60 percent in each of the three ASM bands (1.3–3, 3–5, and 5–12 keV). The detected period is remarkably close to what was suggested for the binary orbital period of XTE J1550–564 (Soria and Wu, IAUC 7184), based on a theoretical relationship between the orbital period and the mean density of the secondary star, which is assumed to have filled its Roche lobe. Considering that few blackhole candidates have been observed to show any x-ray modulations at the binary orbital period, the detection of such modulation in XTE J1550–564 perhaps implies a special environment or geometry of the binary system, like in the well-known case of Cyg X-1 (Wen et al. 1999, Ap.J., in press). Future optical observations of the source in x-ray quiescence can hopefully shed more light on the issue.”

SUPERNOVA 1999cl in NGC 4501 G. Masi, Ceccano, Italy, reports the following precise position for SN

1999cl: α = 12^{h}31^{m}56^{s}.01, δ = +14^{o}25′35′′.3 (equinox 2000.0); on June 5.858 UT, he obtained an unfiltered CCD magnitude of 14.1. D. Hanžl, Brno, Czech Republic, reports the following CCD magnitude: June 2.914, R = 14.7. Visual magnitude estimates: June 3.881, 14.6 (K. Hornoch, Lelekovice, Czech Republic); 4.96, 14.2 (J. Carvajal, Madrid, Spain); 5.883, 14.4 (Hornoch).

1999 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green